OpenStudy (hdrager):

Write the equation of a sine function that has the given characteristics. Please show the process for me if possible. Amplitude: 4 Period: 3pi

jimthompson5910 (jim_thompson5910):

The general sine function is this \[\Large y = A\sin(Bx - C) + D\] |A| = amplitude T = 2pi/B is the period C/B is the phase shift D = used to find the midline In this case, amplitude = 4 so A = 4 period = 3pi so T = 3pi. Use this to find the value of B C = 0 and D = 0

OpenStudy (hdrager):

@jim_thompson5910 What is T?

OpenStudy (hdrager):

And I;m mostly stuck on finding B when forming the equation out of the given info

jimthompson5910 (jim_thompson5910):

T is the period T is used instead of P. I'm not 100% sure why but I'm assuming T refers to time or something

OpenStudy (hdrager):

ah okay

jimthompson5910 (jim_thompson5910):

T = 3pi T = 2pi/B 3pi = 2pi/B 3 = 2/B B = ???

OpenStudy (hdrager):

Ohhhhh okay its 2/3

jimthompson5910 (jim_thompson5910):

yep so the trig equation would be... \[\Large y = A\sin(Bx - C) + D\] \[\Large y = 4\sin\left(\frac{2}{3}x - 0\right) + 0\] \[\Large y = 4\sin\left(\frac{2}{3}x\right)\]

OpenStudy (hdrager):

Sweet thanks so much

jimthompson5910 (jim_thompson5910):

no problem